OpenStudy (anonymous):

x = {3/2} x = {0} x = {3/2, 11/2} No Solution

OpenStudy (anonymous):



OpenStudy (anonymous):

@zordoloom

OpenStudy (anonymous):

@derrel1959 @phi

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

@zordoloom yeah.

OpenStudy (anonymous):

x = {3/2, 11/2}

OpenStudy (anonymous):

Solving for x 7 + -2x = 4 Move all terms containing x to the left, all other terms to the right. Add '-7' to each side of the equation. 7 + -7 + -2x = 4 + -7 Combine like terms: 7 + -7 = 0 0 + -2x = 4 + -7 -2x = 4 + -7 Combine like terms: 4 + -7 = -3 -2x = -3 Divide each side by '-2'. x = 1.5